OpenStudy (anonymous):

Which function below has the largest slope? g(x) g(x)=10x - 4 h(x) Lance gains 1 level each time he plays his favorite video game. He started at level 10.

OpenStudy (anonymous):

f(x) g(x) h(x) The functions all have the same slope

OpenStudy (anonymous):

H

OpenStudy (anonymous):

@Blondie_Forreal how come?

ganeshie8 (ganeshie8):

where is f(x) ?

OpenStudy (anonymous):

f(x) x f(x) -2 4 4 12 6 20

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

ty :)

ganeshie8 (ganeshie8):

lets find slope of each function , so we can knw which one is greater

ganeshie8 (ganeshie8):

slope of f(x) = ? u knw how to get slope ?

OpenStudy (anonymous):

no math isnt my subject at all /:

ganeshie8 (ganeshie8):

lol ok :) slope is easy to calculate when u have table slope = change in y / change in x

ganeshie8 (ganeshie8):

x f(x) -2 4 4 12 <<< this one 6 20 <<< and this one

ganeshie8 (ganeshie8):

20 - 12 ------ 6 - 4

ganeshie8 (ganeshie8):

8 --- 2

ganeshie8 (ganeshie8):

4

ganeshie8 (ganeshie8):

so. slope of f(x) = 4

ganeshie8 (ganeshie8):

lets find slope of g(x)

ganeshie8 (ganeshie8):

g(x)=10x - 4

OpenStudy (anonymous):

-40?

ganeshie8 (ganeshie8):

u knw how to find slope, when equation is given ha ? :)

ganeshie8 (ganeshie8):

noo

ganeshie8 (ganeshie8):

when equation is like below, y = mx + b m is the slope

ganeshie8 (ganeshie8):

g(x)=10x - 4 ^

ganeshie8 (ganeshie8):

so, slope of g(x) is 10

OpenStudy (anonymous):

so the answer would have to be (g)x?

ganeshie8 (ganeshie8):

yup !!

OpenStudy (anonymous):

okay thankyou!

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

@ganeshie8 i need help with one more question please..

OpenStudy (anonymous):

this one is smallest slope

ganeshie8 (ganeshie8):

sure, post the q

OpenStudy (anonymous):

Which function below has the smallest slope? f(x) f(x) = 5x - 4 g(x) graph going through (0, 1) and (2, 5)

OpenStudy (anonymous):

h(x) x h(x) 1 14 4 15 7 16

OpenStudy (anonymous):

wrong question hold on give me one second

ganeshie8 (ganeshie8):

okiee

OpenStudy (anonymous):

Its this one ... Which function below has the smallest slope? f(x) f(x) = 5x - 4 h(x) x h(x) 1 14 4 15 7 16

ganeshie8 (ganeshie8):

oly two functions they gave us ? f(x), h(x)

OpenStudy (anonymous):

yeaa im guessing

OpenStudy (anonymous):

thats all the problem stated then the multiple choice answers

ganeshie8 (ganeshie8):

good :) less work for us

ganeshie8 (ganeshie8):

when u see an equation like below :- f(x) = 5x - 4 ^

ganeshie8 (ganeshie8):

blindly take that, slope = 5 (wat ever sticking to x)

ganeshie8 (ganeshie8):

so, slope of f(x) = 5

OpenStudy (anonymous):

so the answer is f(x)?

ganeshie8 (ganeshie8):

x h(x) 1 14 4 15 7 16 slope = 15-14 / 4-1 = 1/3

ganeshie8 (ganeshie8):

slope of h(x) = 1/3

ganeshie8 (ganeshie8):

so... h(x) has the SMALLEST slope

ganeshie8 (ganeshie8):

read thte q again, we're supposed to find the funciton that has smallest slope.

OpenStudy (anonymous):

I was thinking of the first one thankyou !! :]

ganeshie8 (ganeshie8):

np :) answer is h(x)